OpenStudy (anonymous):

if f(x) =x^2-x^2-5, find f(-2/3)

jimthompson5910 (jim_thompson5910):

is it really x^2 - x^2 - 5 ? or is there a typo?

OpenStudy (anonymous):

x^3-x^2-5

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

f(x) = x^3-x^2-5 f(-2/3) = (-2/3)^3-(-2/3)^2-5 ... replace ALL copies of x with -2/3 f(-2/3) = -8/27-4/9-5 I'll let you finish up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-155/7

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

oh sorry, it should be -155/27

jimthompson5910 (jim_thompson5910):

missed the 2

OpenStudy (anonymous):

I have seen where I made the mistake....i did -8/27-(-4/9)-7

jimthompson5910 (jim_thompson5910):

i see

OpenStudy (anonymous):

oh yea..sorry for that

jimthompson5910 (jim_thompson5910):

no worries

OpenStudy (anonymous):

you still have time, I got few more questions.

jimthompson5910 (jim_thompson5910):

I can do one more

OpenStudy (anonymous):

ok